OpenStudy (anonymous):

look at the following points. (4,0), (3,-1), (6,3), (2,-4) which are solutions to y=x-4 choose all the correct answers. -6,3 -4,0 -3,-1 -2, -4

OpenStudy (anonymous):

the points in the question don't match the options.

OpenStudy (anonymous):

well that's what i thought im so confused

OpenStudy (anonymous):

Maybe it was a mistake :/ The answers all have opposite x-coordinates, and none of them work. Maybe work it with the numbers at the top and choose the closest point

OpenStudy (anonymous):

like to test (4, 0) plug in 4 for x and 0 for y. y = x -4 0 = 4 - 4 0 = 0 this is true so (4, 0) works. So maybe select (-4, 0) as one of the answers. The question has an obvious mistake, so take a screenshot and show it to your teacher if you get it wrong.